OpenStudy (tsunderemailbox):

Determine which ordered pair is not a solution of y = –5x – 4, Explain? (10, –52) (7, –39) (–7, 31) (8, –44)

Mehek (mehek14):

Plug in the x value and y value Start with A x = 10 , y = -52 -52 = -5 * 10 - 4 Does this come out as true?

OpenStudy (tsunderemailbox):

No, That would be -54, Right?

Mehek (mehek14):

Right so it is `NOT` a solution.

OpenStudy (tsunderemailbox):

Okay, So A is out of the question now

Mehek (mehek14):

The question said to look for the ordered pair that is `not` the solution. :P

OpenStudy (tsunderemailbox):

Oh, My bad. >_<, Sorry. Really stressed out with this end of year work

OpenStudy (tsunderemailbox):

Thank you for your help. ^_^!

Mehek (mehek14):

You can check with B, C, and D -39 = -5 * 7 - 4 31 = -5 * - 7 - 4 -44 = -5 * 8 - 4 It's ok :) and you're welcome ^_^